system of linear equations

Mary has 40 bills in her purse, consisting of 20-peso bills and 50-bills. The total amount of the bills is ₱ 1,700. How many of each kind of bill are there in her purse?
Answer #1

Given: Mary has 40 bills in  total. and total amount in wallet is 1700 peso. \(\rightarrow\) (1)

Also, said that bills are 20 peso and 50 peso bills. \(\rightarrow\)(2)

Let ' \(x\) ' be the number of bills of 20 peso

So, \(\quad 50\) peso bils \(=(40-x)\)

$$ \begin{array}{l} \text { By (1) and (2) } \\ 20 x+(40-x) \times 50=1700 \\ 20 x+(40 \times 50)-50 x=1700 \\ x=10 \end{array} $$

Number of bills of 20 peso \(=10\)

\(\begin{aligned} \text { Number of bills of } 50 \text { peso } &=(40-x) \\ &=(40-10)=30 \end{aligned}\)

answered by: Insouse